OpenStudy (anonymous):

find the equation of the line tangent to the function at the given point. f(x)=5x^2 at x=10

OpenStudy (anonymous):

Find the slope: f'(x)=10x f'(x)= 10*10 = 100 Solve for y: \[f(x) = 5x ^{2} = 5 * 100 = 500\] y = 100x-500